# Planet Migration in Protoplanetary Disks with Rims

### v1

@@ 1 Introduction
Protoplanetary disks (PPDs) are residuals of star formation and provide the initial environment for planetary assembly, setting the stage for the demographics of mature planetary systems (Hartmann2001). Planetary formation and evolution are intrinsically linked to tidal interactions with the protoplanetary disk (ida2004; Armitage2011). Low-mass planets, including super-Earths and Neptune-mass planets, usually exert only a modest perturbation on their disk’s structure. The net torque on low mass planets from the disk drives rapid type I migration (goldreich1980; tanaka2002; paardekooper2011; Lubow2010; tanaka2002; McNally2019; Ward1997). Massive planets strongly perturb their disk and open deep gaps in the gaseous disk through gravitational torques exerted at Lindblad and co-rotation resonances (lin1986a; Calvet2002), facilitating angular momentum transport in a viscous-like manner (lin1986b; lin1993). This is the type II regime (Nelson2000; Lubow2006; Drmann2015; Dong2015a; Kanagawa2018; Robert2018). Both processes can drive the efficient inward migration of the embedded planet, which could explain the origin of close-in planets (lin1996).
@@
Recent hydrodynamical simulations indicate that migration rates in the type I and type II regimes can be described by a more generalized migration prescription after considering the gap-opening effect (Kanagawa2018). The migration rate from these simulations, even after considering the suppression due to the gap-opening effect, still poses a challenge to the retention of cold Jupiters when compared with the observed period distribution (Nelson2000; Ida2013). Many scenarios have been explored to identify processes which may lead to the slowing down of migration speed or even the reversal of migration direction (e.g., Masset2006; Lega2014; Chen2020; Dempsey2021; li2024; Wu2024; Laune2024; Hou2024; Wu2025; Liu2025).
@@
High-resolution ALMA observations of the protoplanetary disk have revealed a series of concentric ring- and gap-like substructures (bright rings and narrow dark annuli/gaps) in the dust continuum emission (e.g., Andrews2016; Andrews2018; Long2018; Segura2020; Guidi2022; Booth2024; Carvalho2024; Teague2025; Sierra2025; Gardner2025). Such kind of sub-structures in the dust continuum could be attributed to the intrinsic consequences of angular momentum transfer mechanisms like the dead zone (Bai2013; Lyra2008; Li2019), magnetohydrodynamics zonal flows (Bai2014; Johansen2009; Krapp2018; Cui2021), magnetized disk winds (Suriano2017; Riols2019), vertical shear instability (Nelson2013; Lin2015), the eccentric cooling instability (Lubow1991; LiJ2021), secular gravitational instability (Youdin2011; Tominaga2019), or induced by the embedded planets (lin1986b; Dong2015a; Zhang2018; Li2019b; Paardekooper2023). Some kinematic evidence for protoplanets forming in ringed-structure disks has emerged from the detection of localized deviations from Keplerian velocity in HD 163296 (Pinte2018; Teague2018).
@@
Most studies on planet-disk interaction have focused on disks with smooth initial surface density profiles. As the annular structures such as rings and gaps could coexist in PPDs, the migration pattern in these disks could be important to shape the architectures of exoplanet system, which are, however, remains relatively unexplored. Masset2006 found that a reversed gas surface density gradient, for instance near a cavity, can generate a large, positive corotation torque for low-mass planets, effectively halting their inward migration. This study establishes a new framework for investigating planetary reverse migration and migration arrest (see also, Liu2025).
@@
The rest of the paper is organized as follows. In the Methods section (§2), we describe the numerical setup for the hydrodynamical simulations. We show the simulation results and implications in §3. The conclusion of this work is presented in §4.
2 Methods
In this section, we describe the simulation setup and diagnostics for the migration of the planet with a disk rim. To investigate the influence of the disk on planetary migration, we adopt the conventional $\alpha$-prescription for steady-state accretion disk models (ShakuraSunyaev1973) and modify the $\alpha$ profile to alter the disk’s surface density distribution. A surface density bump can be achieved by implementing a dead zone in the $\alpha$ profile, while a surface density gap can be produced by establishing an $\alpha$ bump configuration.
2.1 Simulation Setup
We globally simulate the hydrodynamical evolution of a planet embedded in a PPD with Athena++ (Stone2020). The continuity equations for the mass and momentum of the fluid are numerically solved in a 2D thin disk by including the gravitational interaction between the planet and the disk. The coordinate of the system is located at the position of the central star with its mass of $M_{*}$. For the 2D disk model, we use the cylindrical coordinate $(r,\ \phi)$ for the simulations.
We follow the same method as in previous works to initialize the disk (e.g., Li2019; Liu2025). The disk surrounding a pre-main sequence (PMS) star of mass $M_{\star}=1.0M_{\odot}$ is initialized with a gas surface density profile following a power-law distribution, which is described as
@@ 2.1 Simulation Setup
2.1.1 The Disk Rim Model
We adopt a transition of the disk viscosity (i.e., the dead zone) to approximate the disk rim model. The dead zone in PPDs is closely related to the magneto-rotational instability (MRI) (Balbus1991), which was long thought to govern accretion in protoplanetary disks. However, low ionization rates in these disks are expected to suppress the MRI within a region of $\sim$1–10 AU, leading to significantly reduced turbulence and angular momentum transport compared to theoretical expectations (Lehmann2022; Gressel2015), which is known as the “dead zone” (Gammie1996; Turner2009; Armitage2011; Turner2014). The formation of this dead zone is attributed to a low ionization fraction, which prevents the gas from coupling effectively with the magnetic field, a dependency that is further influenced by the dust properties within the disk (Dzyurkevich2013). The specific location and radial extent of the dead zone are highly uncertain, as they are governed by complex factors including the sources of ionization, the disk’s physical structure, and the local magnetic field strength.
The dead zone typically manifests as a sharp viscosity gradient at a distance of a few astronomical units from the central star (Armitage2011). In fact, recent state-of-the-art magnetohydrodynamical models show that the inner parts of protoplanetary disks (between $\sim$ 1-20 AU) are largely laminar, with angular momentum transport driven by magneto-thermal winds and laminar Maxwell stresses (Gressel2015; Bai2015).
The gas viscosity is numerically modeled as $\alpha$-prescription with $\nu=\alpha H^{2}\Omega$ (ShakuraSunyaev1973), where $\Omega$ is the angular velocity. To implement the dead zone in the disk, we adopt an $\alpha$ profile as (Pinilla2016)
@@ 2.1.1 The Disk Rim Model
The configuration for the disk viscosity profile has a gap in the $\alpha$ profile with $\alpha=\alpha_{\rm dead}$ between $r_{\rm dz\_in}$ and $r_{\rm dz\_out}$, while it remains constant for the background disk. The parameter $\delta$ controls the transition slope of the viscosity between the dead zone and the active zone. This viscosity profile can mimic the coupled effect of the dead zone and the MHD wind (Pinilla2016). For our Jupiter-mass planet, $\alpha_{\rm dead}=10^{-4}$ to amplify the effect of the dead zone, but $\alpha_{\rm dead}=10^{-3}$ for the super-Earth mass planet. We have also tested $\alpha_{\rm dead}=10^{-3}$ for the Jupiter-mass planet, which does not change our results significantly. For both cases, the background viscosity parameter is set to $\alpha_{\rm active}=10^{-2}$. The conceptual graph is shown in Fig. 1(a). Although the $\alpha$-profile is hypothesized in this work, this is simply used to produce the disk rim and is well motivated by numerical simulations and high-resolution observations, as mentioned above. Different models generating the same disk rim should not change our main results.
Figure 1: Schematic of dead zone structure (panel a), planetary migration directions (panel b), and torque distribution relative to dead zone boundaries (panels c and d). In panel (a) is a conceptual graph of the disk and dead zone. The gray region in the figure denotes the dead zone (with initial $\alpha_{\rm dead}=10^{-4}$ for Jupiter and $\alpha_{\rm dead}=10^{-3}$ for super-Earth) , while the remaining disk regions have initial $\alpha_{\rm active}=10^{-2}$.
@@
2.1.2 Planet-Disk Interaction
@@
where $\bm{r}_{\rm p}$ denotes the position vector of the planet and $\epsilon$ is the softening length. The second term on the right-hand side represents the indirect term arising from the use of a heliocentric coordinate system. We have considered two different planet masses. One is for the Jupiter case with the planet mass set to $M_{\rm p}=0.001\ M_{\odot}$, another is for the super-Earth case with its mass of $M_{\rm p}=3\times 10^{-5}\ M_{\odot}$. The softening length is set as $\epsilon=0.6\ H$ for the super-Earth case, and a larger softening length $\epsilon=0.8\ H$ is set for the Jupiter case to smooth out the fluctuated torque around the planet. Both of them are very close to the calibrated softening length from the 3D simulations (Muller2012; Cordwell2025). The planet is fixed in a circular orbit at a distance $r_{\rm p}=r_{0}$ with $\Omega_{\rm p}=\Omega_{0}$, where $\Omega_{0}$ is the Keplerian frequency at $r_{0}$.
We also explore the effect of accretion on the planet migration with a dead zone. For the active accretion of an embedded planet, we adopt a sinkhole approach around the planet, following previous studies (Li2023; li2024). The accretion process is governed by the sinkhole radius $r_{a}$ and a mass removal rate $\eta$, expressed in units of the local Keplerian frequency $\Omega_{0}$. Within each numerical time step, a uniform fraction of mass is removed from every grid cell located within a distance $\delta r<r_{a}$ (where $\delta r=|\bm{r}-\bm{r}_{\rm p}|$) from the accretor. This procedure ensures that once the surface density profile inside the sinkhole reaches a steady state, the mass removal rate converges to the integrated mass flux into the sinkhole. In our accreting planet setup, we use the parameters $r_{a}=0.1\ r_{\rm H}$, $\eta=10\Omega_{0}$ where $r_{\rm H}=r_{\rm p}(M_{\rm p}/3M_{*})^{1/3}$ is the Hill radius, and the softening length is the same as the accretion radius.
@@
2.2 Data Analysis
2.2.1 Total Torque Calculation
@@
2.2.2 Type I Migration Torque
@@
The gravitational potential of the planet can be expanded in a Fourier series as (Kley2012):
@@
where the expression for the coefficient $\psi_{m}$ is given by (Meyer1987).
@@
The Lindblad torque in each $m$ mode is given by (Chen2020),
@@
The expression for the corotation torque is given by (goldreich1980),
@@
Figure 3: Radial surface density (averaged) profiles and radial torque (sum) profiles for the models with a Jupiter mass planet. Different panels correspond to different runs (enclosed by the purple ‘L’ shaped structures) as shown in the panel(c) of Fig. 1. The simulation parameters are set with $M_{\mathrm{p}}=M_{\mathrm{J}}$. The inner and outer dead-zone boundaries $r_{\mathrm{dz\_in}}$ and $r_{\mathrm{dz\_out}}$ are indicated in each panel. For the left column (top to bottom): the top panel employs $r_{\mathrm{dz\_in}}=0.8\,r_{0}$, $r_{\mathrm{dz\_out}}=1.3\,r_{0}$; the middle panel uses $r_{\mathrm{dz\_in}}=0.8\,r_{0}$, $r_{\mathrm{dz\_out}}=1.4\,r_{0}$; and the bottom panel is specified as $r_{\mathrm{dz\_in}}=0.7\,r_{0}$, $r_{\mathrm{dz\_out}}=1.4\,r_{0}$. For the right column (top to bottom): the top panel adopts $r_{\mathrm{dz\_in}}=0.7\,r_{0}$, $r_{\mathrm{dz\_out}}=1.2\,r_{0}$; the middle panel has $r_{\mathrm{dz\_in}}=0.6\,r_{0}$, $r_{\mathrm{dz\_out}}=1.2\,r_{0}$; and the bottom panel corresponds to $r_{\mathrm{dz\_in}}=0.6\,r_{0}$, $r_{\mathrm{dz\_out}}=1.3\,r_{0}$. The red solid curves trace the simulated surface density distribution $\Sigma(r)$, while orange solid lines indicate the initial surface density slope $\Sigma_{0}(r)$ for reference. The dead zone is marked by gray dotted lines. The blue solid curve represents the radial cumulative torque, which is azimuthally integrated.
3 Results

### v3

@@ I Introduction
Protoplanetary disks (PPDs) are residuals of star formation and provide the initial environment for planetary assembly, setting the stage for the demographics of mature planetary systems (Hartmann, 2001). Planetary formation and evolution are intrinsically linked to tidal interactions with the protoplanetary disk (Ida and Lin, 2004; Armitage, 2011). Low-mass planets, including super-Earths and Neptune-mass planets, usually exert only a modest perturbation on their disk’s structure. The net torque on low mass planets from the disk drives rapid type I migration (Goldreich and Tremaine, 1980; Tanaka et al., 2002; Paardekooper et al., 2011; Lubow and Ida, 2010; Tanaka et al., 2002; McNally et al., 2019; Ward, 1997). Massive planets strongly perturb their disk and open deep gaps in the gaseous disk through gravitational torques exerted at Lindblad and co-rotation resonances (Lin and Papaloizou, 1986a; Calvet et al., 2002), facilitating angular momentum transport in a viscous-like manner (Lin and Papaloizou, 1986b, 1993). This is the type II regime (Nelson et al., 2000; Lubow and D’Angelo, 2006; Dürmann and Kley, 2015; Dong et al., 2015; Kanagawa et al., 2018; Robert et al., 2018). Both processes can drive the efficient inward migration of the embedded planet, which could explain the origin of close-in planets (Lin et al., 1996).
@@
Recent hydrodynamical simulations indicate that migration rates in the type I and type II regimes can be described by a more generalized migration prescription after considering the gap-opening effect (Kanagawa et al., 2018). The migration rate from these simulations, even after considering the suppression due to the gap-opening effect, still poses a challenge to the retention of cold Jupiters when compared with the observed period distribution (Nelson et al., 2000; Ida et al., 2013). Many scenarios have been explored to identify processes which may lead to the slowing down of migration speed or even the reversal of migration direction (e.g., Masset et al., 2006; Lega et al., 2014; Chen et al., 2020; Dempsey et al., 2021; Li et al., 2024; Wu and Li, 2024; Laune et al., 2024; Hou and Yu, 2024; Wu and Chen, 2025; Liu et al., 2025; Pan et al., 2026; Ida et al., 2026).
@@
High-resolution ALMA observations of the protoplanetary disk have revealed a series of concentric ring- and gap-like substructures (bright rings and narrow dark annuli/gaps) in the dust continuum emission (e.g., Andrews et al., 2016, 2018; Long et al., 2018; Segura-Cox et al., 2020; Guidi et al., 2022; Booth et al., 2024; Carvalho et al., 2024; Teague et al., 2025; Sierra et al., 2025; Gardner et al., 2025). Such kind of sub-structures in the dust continuum could be attributed to the intrinsic consequences of angular momentum transfer mechanisms like the dead zone (Bai and Stone, 2013; Lyra et al., 2008; Li et al., 2019b), magnetohydrodynamics zonal flows (Bai and Stone, 2014; Johansen et al., 2009; Krapp et al., 2018; Cui and Bai, 2021), magnetized disk winds (Suriano et al., 2017; Riols and Lesur, 2019), vertical shear instability (Nelson et al., 2013; Lin and Youdin, 2015), the eccentric cooling instability (Lubow, 1991; Li et al., 2021), secular gravitational instability (Youdin, 2011; Tominaga et al., 2019), or induced by the embedded planets (Lin and Papaloizou, 1986b; Dong et al., 2015; Zhang et al., 2018; Li et al., 2019a; Paardekooper et al., 2023). Some kinematic evidence for protoplanets forming in ringed-structure disks has emerged from the detection of localized deviations from Keplerian velocity in HD 163296 (Pinte et al., 2018; Teague et al., 2018).
@@
Most studies on planet-disk interaction have focused on disks with smooth initial surface density profiles. As the annular structures such as rings and gaps could coexist in PPDs, the migration pattern in these disks could be important to shape the architectures of exoplanet system, which are, however, remains relatively unexplored. Masset et al. (2006) found that a reversed gas surface density gradient, for instance near a cavity, can generate a large, positive corotation torque for low-mass planets, effectively halting their inward migration. This study establishes a new framework for investigating planetary reverse migration and migration arrest (see also, Liu et al., 2025).
@@
The rest of the paper is organized as follows. In the Methods section (§II), we describe the numerical setup for the hydrodynamical simulations. We show the simulation results and implications in §III. The conclusion of this work is presented in §IV.
II Methods
In this section, we describe the simulation setup and diagnostics for the migration of the planet with a disk rim. To investigate the influence of the disk on planetary migration, we adopt the conventional $\alpha$-prescription for steady-state accretion disk models (Shakura and Sunyaev, 1973) and modify the $\alpha$ profile to alter the disk’s surface density distribution. A surface density bump can be achieved by implementing a dead zone in the $\alpha$ profile, while a surface density gap can be produced by establishing an $\alpha$ bump configuration.
II.1 Simulation Setup
We globally simulate the hydrodynamical evolution of a planet embedded in a PPD with Athena++ (Stone et al., 2020). The continuity equations for the mass and momentum of the fluid are numerically solved in a 2D thin disk by including the gravitational interaction between the planet and the disk. The coordinate of the system is located at the position of the central star with its mass of $M_{*}$. For the 2D disk model, we use the cylindrical coordinate $(r,\ \phi)$ for the simulations.
We follow the same method as in previous works to initialize the disk (e.g., Li et al., 2019b; Liu et al., 2025). The disk surrounding a pre-main sequence (PMS) star of mass $M_{\star}=1.0M_{\odot}$ is initialized with a gas surface density profile following a power-law distribution, which is described as
@@ II.1 Simulation Setup
II.1.1 The Disk Rim Model
We adopt a transition of the disk viscosity (i.e., the dead zone) to approximate the disk rim model. The dead zone in PPDs is closely related to the magneto-rotational instability (MRI) (Balbus and Hawley, 1991), which was long thought to govern accretion in protoplanetary disks. However, low ionization rates in these disks are expected to suppress the MRI within a region of $\sim$1–10 AU, leading to significantly reduced turbulence and angular momentum transport compared to theoretical expectations (Lehmann and Lin, 2022; Gressel and Pessah, 2015), which is known as the “dead zone” (Gammie, 1996; Turner and Drake, 2009; Armitage, 2011; Turner et al., 2014). The formation of this dead zone is attributed to a low ionization fraction, which prevents the gas from coupling effectively with the magnetic field, a dependency that is further influenced by the dust properties within the disk (Dzyurkevich et al., 2013). The specific location and radial extent of the dead zone are highly uncertain, as they are governed by complex factors including the sources of ionization, the disk’s physical structure, and the local magnetic field strength.
The dead zone typically manifests as a sharp viscosity gradient at a distance of a few astronomical units from the central star (Armitage, 2011). In fact, recent state-of-the-art magnetohydrodynamical models show that the inner parts of protoplanetary disks (between $\sim$ 1-20 AU) are largely laminar, with angular momentum transport driven by magneto-thermal winds and laminar Maxwell stresses (Gressel and Pessah, 2015; Bai, 2015).
The gas viscosity is numerically modeled as $\alpha$-prescription with $\nu=\alpha H^{2}\Omega$ (Shakura and Sunyaev, 1973), where $\Omega$ is the angular velocity. To implement the dead zone in the disk, we adopt an $\alpha$ profile as (Pinilla et al., 2016)
@@ II.1.1 The Disk Rim Model
The configuration for the disk viscosity profile has a gap in the $\alpha$ profile with $\alpha=\alpha_{\rm dead}$ between $r_{\rm dz\_in}$ and $r_{\rm dz\_out}$, while it remains constant for the background disk. The parameter $\delta$ controls the transition slope of the viscosity between the dead zone and the active zone. This viscosity profile can mimic the coupled effect of the dead zone and the MHD wind (Pinilla et al., 2016). For our Jupiter-mass planet, $\alpha_{\rm dead}=10^{-4}$ to amplify the effect of the dead zone, but $\alpha_{\rm dead}=10^{-3}$ for the super-Earth mass planet. We have also tested $\alpha_{\rm dead}=10^{-3}$ for the Jupiter-mass planet, which does not change our results significantly. For both cases, the background viscosity parameter is set to $\alpha_{\rm active}=10^{-2}$. The conceptual graph is shown in Fig. 1(a). Although the $\alpha$-profile is hypothesized in this work, this is simply used to produce the disk rim and is well motivated by numerical simulations and high-resolution observations, as mentioned above. Different models generating the same disk rim should not change our main results (see also discussions in Section III.6).
Figure 1: Schematic of dead zone structure (panel a), planetary migration directions (panel b), and torque distribution relative to dead zone boundaries (panels c and d). Panel a is a conceptual graph of the disk and dead zone. The gray region in the figure denotes the dead zone (with initial $\alpha_{\rm dead}=10^{-4}$ for Jupiter and $\alpha_{\rm dead}=10^{-3}$ for super-Earth) , while the remaining disk regions have initial $\alpha_{\rm active}=10^{-2}$. Panel b illustrates the migration directions of Jupiters and super-Earths in a protoplanetary disk with a surface density bump. The solid circle denotes Jupiter, while the symbol $\bigoplus"$ denotes the super-Earth. Blue indicates inward migration, and red indicates outward migration. In panels c and d, the x-axis represents $\Delta_{\rm dz\_in}$, defined as the distance from the planet to the inner boundary of the dead zone, and the y-axis represents $\Delta_{\rm dz\_out}$, defined as the distance from the planet to the outer boundary of the dead zone. Symbols enclosed by numbers indicate different regions where the planet is located relative to the ring/gap. The color of the points corresponds to the torque value: blue points represent negative torque, corresponding to inward migration, and red points represent positive torque, corresponding to outward migration.
@@
II.1.2 Planet-Disk Interaction
@@
where $\bm{r}_{\rm p}$ denotes the position vector of the planet, $\epsilon$ is the softening length and $q$ represents the mass ratio $M_{p}/M_{\odot}$. The second term on the right-hand side represents the indirect term arising from the use of a heliocentric coordinate system. We have considered two different planet masses. One is for the Jupiter case with the planet mass set to $M_{\rm p}=0.001\ M_{\odot}$, another is for the super-Earth case with its mass of $M_{\rm p}=3\times 10^{-5}\ M_{\odot}$. The softening length is set as $\epsilon=0.6\ H$ for the super-Earth case, and a larger softening length $\epsilon=0.8\ H$ is set for the Jupiter case to smooth out the fluctuated torque around the planet. Both of them are very close to the calibrated softening length from the 3D simulations (Müller et al., 2012; Cordwell et al., 2025). The planet is fixed in a circular orbit at a distance $r_{\rm p}=r_{0}$ with $\Omega_{\rm p}=\Omega_{0}$, where $\Omega_{0}$ is the Keplerian frequency at $r_{0}$.
We also explore the effect of accretion on the planet migration with a dead zone. For the active accretion of an embedded planet, we adopt a sinkhole approach around the planet, following previous studies (Li et al., 2023, 2024). The accretion process is governed by the sinkhole radius $r_{a}$ and a mass removal rate $\eta$, expressed in units of the local Keplerian frequency $\Omega_{0}$. Within each numerical time step, a uniform fraction of mass is removed from every grid cell located within a distance $\delta r<r_{a}$ (where $\delta r=|\bm{r}-\bm{r}_{\rm p}|$) from the accretor. This procedure ensures that once the surface density profile inside the sinkhole reaches a steady state, the mass removal rate converges to the integrated mass flux into the sinkhole. In our accreting planet setup, we use the parameters $r_{a}=0.1\ r_{\rm H}$, $\eta=10\Omega_{0}$ where $r_{\rm H}=r_{\rm p}(M_{\rm p}/3M_{*})^{1/3}$ is the Hill radius, and the softening length is the same as the accretion radius.
@@
II.2 Data Analysis
II.2.1 Total Torque Calculation
@@
II.2.2 Type I Migration Torque
@@
The gravitational potential of the planet can be expanded in a Fourier series as (Kley and Nelson, 2012):
@@
where the expression for the coefficient $\psi_{m}$ is given by (Meyer-Vernet and Sicardy, 1987).
@@
The Lindblad torque in each $m$ mode is given by (Chen et al., 2020),
@@
The expression for the corotation torque is given by (Goldreich and Tremaine, 1980),
@@
Figure 3: Radial surface density (averaged) profiles and radial torque (sum) profiles for the models with a Jupiter mass planet. Different panels correspond to different runs (enclosed by the purple ‘L’ shaped structures) as shown in the panel c of Fig. 1. The simulation parameters are set with $M_{\mathrm{p}}=M_{\mathrm{J}}$. The inner and outer dead-zone boundaries $r_{\mathrm{dz\_in}}$ and $r_{\mathrm{dz\_out}}$ are indicated in each panel. For the left column (top to bottom): the top panel employs $r_{\mathrm{dz\_in}}=0.8\,r_{0}$, $r_{\mathrm{dz\_out}}=1.3\,r_{0}$; the middle panel uses $r_{\mathrm{dz\_in}}=0.8\,r_{0}$, $r_{\mathrm{dz\_out}}=1.4\,r_{0}$; and the bottom panel is specified as $r_{\mathrm{dz\_in}}=0.7\,r_{0}$, $r_{\mathrm{dz\_out}}=1.4\,r_{0}$. For the right column (top to bottom): the top panel adopts $r_{\mathrm{dz\_in}}=0.7\,r_{0}$, $r_{\mathrm{dz\_out}}=1.2\,r_{0}$; the middle panel has $r_{\mathrm{dz\_in}}=0.6\,r_{0}$, $r_{\mathrm{dz\_out}}=1.2\,r_{0}$; and the bottom panel corresponds to $r_{\mathrm{dz\_in}}=0.6\,r_{0}$, $r_{\mathrm{dz\_out}}=1.3\,r_{0}$. The red solid curves trace the simulated surface density distribution $\Sigma(r)$, while gray dashed lines indicate the initial surface density slope $\Sigma_{0}(r)$ for reference. The dead zone is marked by gray dotted lines. The blue solid curve represents the radial cumulative torque, which is azimuthally integrated. The corresponding differential torque density profiles are shown by the purple solid lines in the inset at the upper right corner.
III Results